If a sound pressure is measured at 2Nm^-2, calculate this in dB(SPL)
1
Expert's answer
2012-06-05T12:01:30-0400
If a sound pressure is measured at 2Nm^-2, calculate this in
dB(SPL)
Reference sound pressure p0 = 20 μPa = 2 × 10-5 Pa ≡ 0 dB
2 Pa = 2
/ 0.02 = approx. 100 dB.